OpenStudy (omgthatgirl):

Solve: x^2-2x+1=4 x=? and x=?

OpenStudy (dtan5457):

Subtract the 4 and solve the quadratic. x^2-2x-3=0 Know how to factor these?

OpenStudy (omgthatgirl):

No...

OpenStudy (dtan5457):

Look for two numbers that add up to -2 and multiply to -3.

OpenStudy (anonymous):

well she has a, b and c for the quadratic.. which would give you x=3, and -1. is factoring easier?

OpenStudy (omgthatgirl):

-1+-1=-2

OpenStudy (omgthatgirl):

i don't really know, pretty confusing to me sorry

OpenStudy (jdoe0001):

hint: \(\large { x^2-2x+1=4\implies x^2-2x+1-4=0 \\ \quad \\ \implies \begin{array}{cccllll} x^2&-2x&-3&=0\\ &-3+1&-3\cdot 1 \end{array} }\)